# OneForecast: A Universal Framework for Global and Regional Weather Forecasting

### v1

@@ 3.5 Long-term and Ensemble Forecasts
As shown in Figure 5, we evaluate long-term forecasts with the two best models on Z500 (500hPa Geopotential). Pangu exhibits patch artifacts in 100-day forecasts, while GraphCast experiences error accumulation that degrades the forecasted physical fields, rendering them physically implausible over time. In contrast, OneForecast achieves stable long-term forecast performance, effectively capturing large-scale atmospheric states without the aforementioned issues. These results highlight OneForecast’s superior capability in maintaining accurate and physically consistent predictions over extended forecasting horizons. In Table 2, we show the results for 10 initial conditions (starting from 00:00 UTC 1 January 2020, and the interval between consecutive initial conditions is 100 time steps). The ensemble forecast (ENS) results are averaged from 50 members. Obviously, the forecasting performance is enhanced for each model when uncertainty is incorporated, and OneForecast still achieves the best performance.
@@ 3.6 Ablation Studies
To verify the effectiveness of the proposed method, as shown in Table 3, we conducted detailed ablation experiments. We introduced the following model variants: (1) OneForecast w/o Region Refined Graph, we remove the region refined mesh from the finest mesh and compute the regional metrics. (2) OneForecast w Region Refined Graph, we reserve the region refined mesh. (3) OneForecast R w/o NGG, we remove the neural nested grid (NGG) method in the regional forecasts and only use the regional data to train the model. (4) OneForecast R BF, we remove the NGG in the regional forecasts and only use the boundary forcing method to train the model. The results of the ablation experiment show that removing any module will reduce the forecast performance, further proving the rationality of the components proposed by OneForecast. (5) OneForecast R w NGG, regional forecast model with NGG. (6) OneForecast w/o MSM, we remove the multi-stream messaging module (MSM) and use a traditional MLP-based messaging module. (7) OneForecast, the full version of OneForecast for the global forecasts. For (1) and (2), we only evaluate the region-refined data. For (3), (4), and (5), we only evaluate the specific regional data. For (6) and (7), we evaluate the global data. And these results are based on 4-day forecast results. Experimental results show that whether it is a global or regional forecast task, the lack of any component will degrade the performance of OneForecast, which proves the effectiveness of the proposed method. And as shown in Figure 7, the proposed MSM can better capture of high and low frequency information, which achieves satisfactory results in long-term forecasts and extreme event forecasts, such as typhoons.
@@
OneForecast w/o Region Refined Graph 0.3751 0.8954
OneForecast w Region Refined Graph 0.2736 0.9342
OneForecast R w/o NGG 0.5850 0.8405
OneForecast R BF 0.4473 0.8734
OneForecast R w NGG 0.3168 0.9325
@@ Appendix A Proofs of Theorems
We consider a multi-stream message passing operator that we denote by $\mathbf{G}$. This operator depends on both node features and dynamic gating on edges. We let $\mathbf{f}\in\mathbb{R}^{N}$ be an arbitrary graph signal, and we write its graph Fourier transform as $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$.
Step 1: Integral analogy of gating on a discrete graph. We first recall that on a continuous domain, a gating operator often admits a representation of the form
@@
Appendix B Data Details
B.1 Dataset
In this section, we introduce the dataset used in this paper detailedly. For the global forecasting, we conduct experiments on the weatherbench2 (Rasp et al., 2024) benchmark, a subset of ERA5 reanalysis data (Hersbach et al., 2020). The weatherbench2 benchmark we used is the version of 1.5° resolution (121 × 240), which spans from 1959 to 2020. This subset contains 5 variables (Z, Q, T, U, V) with 13 pressure levels (50 hPa, 100 hPa, 150 hPa, 200 hPa, 250 hPa, 300 hPa, 400 hPa, 500 hPa, 600 hPa, 700 hPa, 850 hPa, 925 hPa and 1,000 hPa) and 4 variables (U10M, V10M, T2M, MSLP) with surface level. For the regional forecasting, a higher resolution data (0.25° resolution) of ERA5 is also used, which can be downloaded from https://cds.climate.copernicus.eu/, the official website of Climate Data Store (CDS). All the data we used are shown in Table 4. For both global and regional forecasts, we use the data from 1959 to 2017 for training, 2018 to 2019 for validating, and 2020 for testing.
@@ B.1 Dataset
B.2 Data preprocessing
Different atmospheric variables have large differences in numerical magnitude. To allow the model focusing on predictions rather than learning the differences between variables, we normalized the data before putting the data into the model. We calculated the mean and standard deviation of all variables using data from 1959 to 2017 (training set). Each variable has a corresponding mean and standard deviation. Before putting the data into the model, we subtracted the change from the respective mean and divided it by the standard deviation.
Appendix C Algorithm
@@
Algorithm 1 OneForecast Framework for Global Weather Forecast
@@
12:     Decoder Project the feature from mesh back to lat-lon grid: ${\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})$, ${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}_{d1}}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime})$, $\mathcal{V}_{f}^{G}={\mathrm{MLP}}_{d2}(\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G})$, $Z_{t+1}=Z_{t}+{\mathcal{V}_{f}^{G}}$
@@
Appendix D Model Details for Global Forecasts
D.1 Earth-specific Region Refined Multi-scale Graph
@@
D.2 Encoder
We first apply embedder MLP to map the data to the latent space, which can be defined as:
@@ D.2 Encoder
where, $x_{embedder}$ is the input of embedder MLP. For the linear function, we set the latent dim to 512. $SiLU(\cdot)$ is the SiLU activation function, $LN(\cdot)$ is the layernorm function. $Z_{t}$, $\mathcal{V}$, $\mathcal{E}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are embedded features of grid nodes, mesh nodes, mesh edges, grid to mesh edges, and mesh to grid edges.
We then project the atmospheric state from the lat-lon grid into the mesh nodes. Specifically, we first update the edge features through an Edge Sum MLP (ESMLP):
$\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}=\mathbf{W}_{e}\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}}$ (24)
$\mathcal{V}_{f}^{G\prime}=\mathbf{W}_{s}\mathcal{V}_{f}^{G},$ (25)
$h^{\prime}=\mathbf{W}_{d}h+\mathbf{b}_{d},$ (26)
$\mathbf{h}_{\mathrm{sum}}=\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}+\mathcal{V}_{f}^{G\prime}+h^{\prime},$ (27)
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\operatorname{LN}\left(\mathbf{W}_{\mathrm{ESMLP}}\sigma\left(h_{out}\right)+\mathbf{b}_{\mathrm{ESMLP}}\right),$ (28)
where, $\mathbf{W}_{e}$, $\mathbf{W}_{s}$, $\mathbf{W}_{d}$ are the linear transformation matrix of grid2mesh edge features, grid node feature, and mesh node features. $\mathbf{W}_{\text{ESMLP }}$ is the linear transformation matrix of output layer. $b_{d}$ is the bias of mesh node during linear transformation, $b_{ESMLP}$ is the bias vector of ESMLP. In summary, the grid2mesh edge update process can be define as:
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}).$ (29)
@@
where, $\sum\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}^{\prime}}$ are the edges that arrives at mesh node.
Then, we update the grid node features using another MLP:
@@
D.3 Multi-stream Messaging
@@
D.4 Decoder
In the decoder, we map the feature from mesh back to lat-lon grids, similar to encoder, we first up the mesh2grid features:
$\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}^{\prime}}=\operatorname{ESMLP}\left(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}}\right).$ (32)
@@ D.4 Decoder
$\mathcal{V}_{f}^{G^{\prime}}=\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G}.$ (34)
@@
$\mathcal{V}_{f}^{G}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G}\right).$ (35)
Appendix E Experiments Details
E.1 Evaluation Metric
We utilize two metrics, RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), to evaluate the forecasting performance, which can be defined as:
@@ E.1 Evaluation Metric
$\small\operatorname{ACC}(k,t)=\sqrt{\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\mathbf{A^{\prime}}_{ij,t}^{k}}{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\right)^{2}\times\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\mathbf{A^{\prime}}_{ij,t}^{k}\right)^{2}}},$ (37)
where $\mathbf{A}_{i,j,t}^{v}$ represents the value of variable v at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from 59 years of training data. To evaluate model performance, RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for variable $v$ a given lead time $\Delta t$.
E.2 Model Training
We train all baseline models and OneForecast using the same training framework. We set the total model training epochs to 110, the initial learning rate is 1e-3, and use the cosine annealing scheduler to adjust the learning rate until the model converged. The model codes of Pangu, Graphcast, and Fengwu we used are released by NVIDIA modulus (https://github.com/NVIDIA/modulus). The model code of Fuxi is obtained by sending an email to the author. For all models, we select the checkpoint that performed best on the validation set for comparative analysis.
E.3 Typhoon Tracking

### v2

@@ 3.5 Long-term and Ensemble Forecasts
As shown in Figure 5, we evaluate long-term forecasts with the two best models on Z500 (500 hPa Geopotential). Pangu exhibits patch artifacts in 100-day forecasts, while GraphCast experiences error accumulation that degrades the forecasted physical fields, rendering them physically implausible over time. In contrast, OneForecast achieves stable long-term forecast performance, effectively capturing large-scale atmospheric states without the aforementioned issues. These results highlight OneForecast’s superior capability in maintaining accurate and physically consistent predictions over extended forecasting horizons. In Table 2, we show the results for 10 initial conditions (starting from 00:00 UTC 1 January 2020, and the interval between consecutive initial conditions is 12 hour). The ensemble forecast (ENS) results are averaged from 50 members. Obviously, in most cases (especially for longer time), the forecasting performance is enhanced for each model when uncertainty is incorporated, and OneForecast still achieves the best performance.
@@ 3.6 Ablation Studies
To verify the effectiveness of the proposed method, as shown in Table 3, we conduct detailed ablation experiments. We introduce the following model variants: (1) OneForecast w/o Region Refined Graph, we remove the region refined mesh from the finest mesh and compute the regional metrics. (2) OneForecast w Region Refined Graph, we reserve the region refined mesh. (3) OneForecast R w/o NNG, we remove the neural nested grid method (NNG) in the regional forecasts and only use the regional data to train the model. (4) OneForecast R BF, we remove the NNG in the regional forecasts and only use the boundary forcing method to train the model. (5) OneForecast R w NNG, regional forecast model with NNG. (6) OneForecast w/o MSM, we remove the multi-stream messaging module (MSM) and use a traditional MLP-based messaging module. (7) OneForecast, the full version of OneForecast for the global forecasts. For (1) and (2), we only evaluate the region-refined data. For (3), (4), and (5), we only evaluate the specific regional data. For (6) and (7), we evaluate the global data. And these results are based on 4-day forecasts for 100 ICs. Experimental results show that whether it is a global or regional forecast task, the lack of any component will degrade the performance of OneForecast, which proves the effectiveness of the proposed method. And as shown in Figure 7, the proposed MSM can better capture of high and low frequency information, which achieves satisfactory results in long-term forecasts and extreme event forecasts, such as typhoons.
@@
OneForecast w/o Region Refined Graph 0.3793 0.6075
OneForecast w Region Refined Graph 0.2609 0.8099
OneForecast R w/o NNG 0.5828 0.2450
OneForecast R BF 0.4428 0.4711
OneForecast R w NNG 0.2180 0.8856
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here. In addition, we propose OneForecast, a global-regional weather forecasting model based on graph neural networks. In the future, we will further study the combination of graph neural networks and dynamic systems to improve the accuracy of forecasts, especially for extreme event forecasts.
@@ Appendix A Proofs of Theorems
We consider a multi-stream message passing operator that we denote by $\mathbf{G}$. This operator depends on both node features and dynamic gating on edges. We let $\mathbf{f}\in\mathbb{R}^{N}$ be an arbitrary graph signal, and we write its graph Fourier transform as $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$. Step 1: Integral analogy of gating on a discrete graph. We first recall that on a continuous domain, a gating operator often admits a representation of the form
@@
Appendix B Related Work
B.1 Deep Learning based Weather Forecasting
B.2 Numerical analysis methods
Appendix C Data Details
C.1 Dataset
In this section, we are going to introduce the dataset we used in this study detailedly. For the global forecasting, we conduct experiments on the WeatherBench2 (Rasp et al., 2024) benchmark, a subset of ERA5 reanalysis data (Hersbach et al., 2020). The WeatherBench2 benchmark we used is the version of 1.5° resolution (121 × 240), which spans from 1959 to 2020. This subset contains 5 variables (Z, Q, T, U, V) with 13 pressure levels (50 hPa, 100 hPa, 150 hPa, 200 hPa, 250 hPa, 300 hPa, 400 hPa, 500 hPa, 600 hPa, 700 hPa, 850 hPa, 925 hPa and 1,000 hPa) and 4 variables (U10M, V10M, T2M, MSLP) with surface level. For the regional forecasting, a higher resolution data (0.25° resolution) of ERA5 is also used, which can be downloaded from https://cds.climate.copernicus.eu/, the official website of Climate Data Store (CDS). All the data we used are shown in Table 4. For both global and regional forecasts, we use the data from 1959 to 2017 for training, 2018 to 2019 for validating, and 2020 for testing.
@@ C.1 Dataset
Table 5: The Params and MACs comparsion of different models.
      Model       Params (M)       MACs (G)
      Pangu (Bi et al., 2023)       23.83       142.39
      Fengwu (Chen et al., 2023a)       153.49       132.83
      Graphcast (Lam et al., 2023)       28.95       1639.26
      Fuxi (Chen et al., 2023b)       128.79       100.96
      OneForecast       24.76       509.27
C.2 Data preprocessing
Different atmosphere and ocean variables have large variations in their magnitude. To allow the model focusing on predictions rather than learning the differences between variables, we normalized the data before putting the data into the model. We calculated the mean and standard deviation of all variables using data from 1959 to 2017 (training set). Each variable has a corresponding mean and standard deviation. Before feeding the data into the model, we first subtract the respective mean and divided it by the standard deviation.
Appendix D Algorithm
@@
Algorithm 1 OneForecast Framework for Global Weather Forecasting
@@
12:     Decoder Project the feature from mesh back to lat-lon grid: ${\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})$, ${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}_{d1}}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime})$, $\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}$, $Z_{t+1}=\mathrm{MLP_{d2}}({\mathcal{V}_{f}^{G}})$
@@
Appendix E Model Details for Global Forecasts
E.1 Earth-specific Region Refined Multi-scale Graph
@@
E.2 Encoder
This paper uses 2 types MLP. We denote the first type as MLP(·), the number of layer is 1, the latent dim is 512, and followed by the SiLU activation function and Layernorm function. And we denote the second type as ESMLP(·), the other hyperparameters are the same as MLP(·), except ESMLP(·) transforms three features (edge features, node features of the corresponding source and destination node) individually through separate linear transformations and then sums them for each edge accordingly. We first apply embedder MLP to map the data to the latent space, which can be defined as:
@@ E.2 Encoder
where, $x_{embedder}$ is the input of embedder MLP. For the linear function, we set the latent dim to 512. $SiLU(\cdot)$ is the SiLU activation function, $LN(\cdot)$ is the layernorm function. $Z_{t}$, $\mathcal{V}$, $\mathcal{E}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are embedded features of grid nodes, mesh nodes, mesh edges, grid to mesh edges, and mesh to grid edges. We then project the atmospheric state from the lat-lon grid into the mesh nodes. Specifically, we first update the edge features through an Edge Sum MLP (ESMLP):
$\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}=\mathbf{W}_{e}\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}},$ (24)
${h_{s}}^{\prime}=\mathbf{W}_{s}h_{s},$ (25)
$h_{d}^{\prime}=\mathbf{W}_{d}h_{d}+\mathbf{b}_{d},$ (26)
$\mathbf{h}_{\mathrm{sum}}=\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}+h_{s}^{\prime}+h_{d}^{\prime},$ (27)
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\operatorname{LN}\left(\mathbf{W}_{\mathrm{ESMLP}}\sigma\left(h_{sum}\right)+\mathbf{b}_{\mathrm{ESMLP}}\right),$ (28)
where, $\mathbf{W}_{e}$, $\mathbf{W}_{s}$, $\mathbf{W}_{d}$ are the linear transformation matrix of grid2mesh edge features, send node feature, and target node features. $\mathbf{W}_{\text{ESMLP }}$ is the linear transformation matrix of output layer. $b_{d}$ is the bias of mesh node during linear transformation, $b_{ESMLP}$ is the bias vector of ESMLP. In summary, the grid2mesh edge update process can be define as:
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f},h_{s},h_{d}).$ (29)
@@
where, $\sum\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}^{\prime}}$ are the edges that arrives at mesh node. Then, we update the grid node features using another MLP:
@@
E.3 Multi-stream Messaging
@@
E.4 Decoder
In the decoder, we map the feature from mesh back to lat-lon grids, similar to encoder, we first update the mesh2grid features:
$\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}^{\prime}}=\operatorname{ESMLP}\left(\mathcal{V}_{f}^{G},h,\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}}\right).$ (32)
@@ E.4 Decoder
$\mathcal{V}_{f}^{G}=\mathcal{V}_{f}^{G}+\mathcal{V}_{f}^{G^{\prime}}.$ (34)
@@
$Z_{t+1}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G}\right).$ (35)
Appendix F Experiments Details
F.1 Evaluation Metric
We utilize four metrics, RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), CSI (Critical Success Index), and SEDI (Symmetric Extremal Dependence Index) to evaluate the forecasting performance, which can be defined as:
@@ F.1 Evaluation Metric
$\small\operatorname{ACC}(k,t)=\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\mathbf{A^{\prime}}_{ij,t}^{k}}{\sqrt{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\right)^{2}\times\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\mathbf{A^{\prime}}_{ij,t}^{k}\right)^{2}}},$ (37)
where $\mathbf{A}_{i,j,t}^{v}$ represents the value of variable v at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from 59 years of training data. To evaluate model performance, RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for variable $k$ at a given lead time $\Delta t$.
$\operatorname{CSI}(k,t)=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}+\mathrm{FN}},$ (38)
$\operatorname{SEDI}(k,t)=\frac{\log(F)-\log(H)-\log(1-F)+\log(1-H)}{\log(F)+\log(H)+\log(1-F)+\log(1-H)},$ (39)
where, true positives (TP) indicate the number of cases in which the state is accurately simulated. False positives (FP) and false negatives (FN) are defined in a similar manner. The false alarm rate is denoted as $F=\frac{\mathrm{FP}}{\mathrm{FP}+\mathrm{TP}}$, while the hit rate is represented as $H=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}$.
F.2 Model Training
For the first type comparison, we train baseline models and OneForecast using the same training framework. We set the total model training epochs to 200, the initial learning rate is 1e-3, and use the cosine annealing scheduler to adjust the learning rate until the model converged. The model codes of Pangu and Graphcast we used are released by NVIDIA modulus (https://github.com/NVIDIA/modulus). The model code of Fuxi is obtained by sending an email to the author. For all models, we select the checkpoint that performed best on the validation set for comparative analysis.
F.3 Typhoon Tracking